OpenStudy (anonymous):

Choose the equation of the line passing through the point (3, -2) and parallel to y = 3x - 6.

OpenStudy (anonymous):

Choose the equation of the line passing through the point (3, -2) and parallel to y = 3x - 6. y = 3x + 7 y = 3x - 11 y = -3x + 11 y = -3x - 12

OpenStudy (lncognlto):

As the lines are parallel, the gradients will be the same. So subst. m = 3 and (3,-2) in y = mx + c. Solve for c and then subst. c and m back in y = mx + c.